Alternative R has a first cost of $82,000, annual M&O costs of $52,000, and a $20,000 salvage value after 5 years. Alternative S has a first cost of $175,000 and a $70,000 salvage value after 5 years, but its annual M&O costs are not known. Determine the M&O costs for alternative S that would yield a required incremental rate of return of 24%.

The M&O cost for alternative S is $_____ .
Answer #1

Let x = M & O costs.

Incremental cash flow analysis.
0 = -93,000 + (-x + 52,000)(P/A,24%,5) + 50,000(P/F,24%,5)

0 = -93,000 + (-x + 52,000)(2.7454) + 50,000(0.3411)

x = $-93000+142760.8+17055/2.7454

x=66815.8/2.7454

x=24337.36
M & O cost for S = $24337.36